# Exploring Behavior-Relevant and Disentangled Neural Dynamics with Generative Diffusion Models

### v1

@@ 5.2 Disentangled Neural Latent Subspace Investigation
Single Latent Factor Behavior Decoding.  In order to verify the disentanglement of the learned neural subspace in BeNeDiff, we evaluate the behavior label decoding performance of each individual latent factor. Specifically, we train a unique linear regressor for each latent factor from the VAE and plot the decoding accuracy as the R-squared value ($R^{2}\%$) in Figure 4. The main observation is that each latent factor is specific to a unique behavior of interest, confirming the orthogonality and clear disentanglement of the inferred latent trajectories from a quantitative perspective.
Neural Observation Signals Reconstruction.  To prevent the VAE from overfitting to the behavioral labels, BeNeDiff also aims to maintain a low reconstruction error for neural activity. We present the quantitative reconstruction results in Table 1, which records the R-squared values ($R^{2}$, in %) and RMSE of the methods. Additionally, we plot the ground-truth neural signals and the reconstructed signals of several methods in a single trial in Figure 5. The main observation is that the neural reconstruction is well-preserved given the behavioral priors. One possible explanation is that the behavioral labels rotate the latent subspace while preserving the necessary information for reconstructing the neural data. The neural signals can be hardly recovered from the behavior labels only. It indicates that the behavior-informed latent does encode significant neural information that is not contained in the behavior labels. Furthermore, we evaluate the disentanglement quality of the latent subspace using the widely-adopted MIG (Mutual Information Gap) metric (Chen et al., 2018), also listed in Table 1. We observe that the learned latent subspace of BeNeDiff significantly enhances disentanglement compared to the vanilla VAE.
Metrics \ Method Session-1 Session-2
Standard VAE Ours Standard VAE Ours
@@ 5.3 Neural Dynamics Exploration of Disentangled Latent Factors
Figure 7: Generated Single-trial Behavioral Videos with Latent Factor Guidance from the bottom view. Compared to baseline methods, we observe that the neural dynamics of latent factor in the results of BeNeDiff show specificity to the “Paw-(y)” movements.
@@
Results Analyses of the “Jaw” Latent Factor from the Side View.  As shown in Figure 6, for the naïve latent manipulation method, the distribution of neural signals often falls outside the original distribution after manipulation, resulting in blurred generated frames. The frame differences are entangled, and the “Jaw” latent factor affects the entire head movement of the mouse, particularly in the first four frames shown. On the other hand, for classifier-free guidance, the generated videos maintain coherent consistency between frames. However, it does not interpret neural dynamics well in this context, resulting in a trajectory with small movements in the “Jaw”. This is because the overall latent trajectory is used as the input to the model and the other latent factors are kept fixed, making it difficult to discriminate the evolution of a single factor effectively. In contrast, the results of BeNeDiff show more specificity to the targeted behavior of interest. The inter-frame differences in BeNeDiff’s results are clearly specified to the “Jaw” movements, and the structure of the neural dynamics is well-preserved and consistent with ground-truth “Jaw” behavior trajectories.
Results Analyses of the “Paw-(y)” Latent Factor from the Bottom View.  As shown in Figure 7, for the naïve latent manipulation method, the generated frames are in a reasonable form. Nevertheless, the frame differences are still intertwined, and the latent factor of “Paw-(y)” heavily affects the “Spout” movement. Meanwhile, for classifier-free guidance, the trajectories focus on the mouse movements, but they are still entangled with the “Chest” movements. In contrast, the results of BeNeDiff show more specificity to the targeted behavior of interest. The inter-frame differences in BeNeDiff’s results are clearly specified to the “Paw-(y)” movements, and the temporal evolution of the neural dynamics is coherent with real-world mouse paw trajectories.
The generated results on all of the six behaviors of interest are in the Appendix and Supplementary Materials.
6 Discussion
Our study introduces BeNeDiff, a novel approach leveraging behavior-informed latent variable models and generative diffusion models to uncover and interpret neural dynamics. Through empirical validation, we demonstrate that BeNeDiff effectively identifies a disentangled neural subspace and synthesizes behavior videos that provide interpretable insights into neural activities associated with distinct behaviors of interest. However, the method relies on behaviors of interest that are disentangled and separate. For more complex dataset and experimental settings, the effectiveness of BeNeDiff needs to be further validated.
@@ Appendix A Methodology details
Training Details of Neural LVM.  Both the probabilistic encoder and decoder of the neural LVM are based on an RNN architecture Fabius and Van Amersfoort [2014]. Mean squared error (MSE) is used for both the neural reconstruction and behavior decoding loss. We use the Adam Optimizer Kingma and Ba [2014] for optimization and the learning rate is set as 0.001. The batch size is uniformly set to 32. The latent subspace factor number is fixed at 6, which is the same as the number of behavior of interest. We employ the dropout technique Srivastava et al. [2014] and the ReLU activation function Rasamoelina et al. [2020] between layers in our probabilistic encoder and decoder neural networks.
Training Details of Video Diffusion Models.  We adopt the architecture of the VDM of 3D-UNet [Ho et al., 2022b] with the $\epsilon$-parameterization. We use both spatial attention and spatial convolutions. The temporal convolutions are used to maintain consistency between frames. The embedding input size to the UNet architecture is set as $32$ and the UNet has three downsampling and upsampling layers. The diffusion timestep is set as $200$. The training batch size is set as $64$, with a learning rate of 0.001.
@@ Appendix B In-depth Investigation on the Neural LVM module across brain regions
Region Metrics SSL CEBRA pi-VAE Ours
[…]
@@ Appendix C Video Generation Results on Various Behaviors of Interests
Figure 8: Generated Single-trial Behavioral Videos with Latent Factor Guidance from the bottom view. Compared to baseline methods, we observe that the neural dynamics of latent factor in the results of BeNeDiff show specificity to the “Paw-(x)” movements.
Figure 9: Generated Single-trial Behavioral Videos with Latent Factor Guidance from the bottom view. Compared to baseline methods, we observe that the neural dynamics of latent factor in the results of BeNeDiff show specificity to the “Spout” movements.
Appendix D Video Generation Results on Various Brain Regions
Figure 10: Learnt Neural Latent Trajectories of BeNeDiff from various brain regions. (A), (B), and (C) represent the latent trajectories for three different trials, respectively. The corresponding generated videos can be found in the supplementary materials.
@@ 8. Answer: [Yes]
  9. Justification: In Section 6 Discussion.

### v3

@@ 5.2 Disentangled Neural Latent Subspace Investigation
We note that we train a unique neural LVM for each individual brain region (single-region), and we evaluate both the behavior decoding and neural reconstruction performance of each brain region-specific neural latent trajectories.
Single Latent Factor Behavior Decoding.  In order to verify the disentanglement of the learned neural subspace in BeNeDiff, we evaluate the behavior label decoding performance of each individual latent factor. Specifically, we train a unique linear regressor for each latent factor from the VAE and plot the decoding accuracy as the R-squared value ($R^{2}\%$). The results of VIS-Right region (the right visual region) are shown in Figure 4. The main observation is that each latent factor is specific to a unique behavior of interest, confirming the orthogonality and clear disentanglement of the inferred latent trajectories from a quantitative perspective.
Neural Observation Signals Reconstruction.  To prevent the VAE from overfitting to the behavioral labels, BeNeDiff also aims to maintain a low reconstruction error for neural activity. Table 1 presents the quantitative reconstruction results compared to baseline methods, including Semi-Supervised Learning (SSL) (Kingma et al., 2014), CEBRA (Schneider et al., 2023), and pi-VAE (Zhou and Wei, 2020). The table records the R-squared values ($R^{2}$, in %) and RMSE for each method. Additionally, we plot the ground-truth neural signals and the reconstructed signals of several methods in a single trial in Figure 5. The main observation is that the neural reconstruction is well-preserved given the behavioral priors. One possible explanation is that the behavioral labels rotate the latent subspace while preserving the necessary information for reconstructing the neural data. The neural signals can be hardly recovered from the behavior labels only. It indicates that the behavior-informed latent does encode significant neural information that is not contained in the behavior labels. Furthermore, we evaluate the disentanglement quality of the latent subspace using the widely-adopted MIG (Mutual Information Gap) metric (Chen et al., 2018), also listed in Table 1. We observe that the learned latent subspace of BeNeDiff significantly enhances disentanglement compared to the vanilla VAE.
Region Metrics SSL CEBRA pi-VAE Ours
[…]
@@ 5.3 Neural Dynamics Exploration of Disentangled Latent Factors
Results Analyses of the Generated Videos.  As shown in Figure 6, for the naïve latent manipulation method, the distribution of neural signals often falls outside the original distribution after manipulation, resulting in blurred generated frames. The frame differences are entangled, and the “Jaw” latent factor affects the entire head movement of the mouse, particularly in the first four frames shown. On the other hand, for classifier-free guidance, the generated videos maintain coherent consistency between frames. However, it does not interpret neural dynamics well in this context, resulting in a trajectory with small movements in the “Jaw”. This is because the overall latent trajectory is used as the input to the model and the other latent factors are kept fixed, making it difficult to discriminate the evolution of a single factor effectively. In contrast, the results of BeNeDiff show more specificity to the targeted behavior of interest. The inter-frame differences in BeNeDiff’s results are clearly specified to the “Jaw” movements, and the structure of the neural dynamics is well-preserved and consistent with ground-truth “Jaw” behavior trajectories. A similar pattern is evident with the other latent factors, as shown in Figures 9, 10, and 11 in the appendix.
5.4 Neural Dynamics Exploration of Disentangled Latent Factors Across Brain Regions
Besides the capability of revealing interpretable neural dynamics of each latent factor associated with behaviors, here we further investigate the neural dynamics differences across brain regions through BeNeDiff. As shown in Figure 7 and Figure 12 in the appendix, we present the 2D neural latent trajectories of two latent factors, specifically related to "Paw-(x)" and "Paw-(y)", across six brain regions for two randomly selected trials. From the starting point of the trial, we observe that the latent trajectories corresponding to the left and right hemispheres of the VIS both show a noticeable change starting earlier. Next, the SSp regions show a large shift in activity, followed by a similar change in the MOs regions. However, it is difficult to clearly visualize the specific motion encoded by each region and to distinguish how different the motions are encoded solely based on neural trajectory plots. This further highlights the need for using a video diffusion model for visualization and interpretation.
In contrast, in the generated behavior video samples of BeNeDiff (as illustrated by the frame differences in Figure 8 and Figure 13 in the appendix), where the "Paw-(x)" and "Paw-(y)" latent factors are activated, the behavioral dynamics encoded by these two latent factors are observed across different brain regions. First, paw movements are detected in the VIS regions before the "Levers" come in. This early activity in VIS could reflect its role in the predictive coding of behaviors, indicating that this region may predict motor movements before they happen. Next, the SSp regions exhibit paw movements that are synchronized with the onset of the "Levers", indicating a potential role for SSp in processing somatosensory feedback. Subsequently, in the MOs regions, paw movements are observed following the "Levers" onset, which is consistent with MOs’ role in motor execution and control, occurring slightly after SSp.
Figure 7: Learnt Neural Latent Trajectories of BeNeDiff across various brain regions. It is difficult to clearly visualize the specific motion encoded by each region and to distinguish how different the motions are encoded across brain regions.
Figure 8: Generated video frame differences across the right hemisphere regions. The red dots in the figure indicate paw appearances.
To sum up, although neural trajectory plots provide a clear temporal sequence of activations across regions, it is challenging to directly visualize the specific behavioral dynamics encoded by each region and to discriminate how they differ. This limitation highlights the necessity for a video diffusion model in BeNeDiff, to better visualize and interpret the encoded behavioral dynamics of each neural latent factor. By synthesizing realistic behavior videos in a generative fashion, BeNeDiff enables us to better understand the unique neural dynamics in each brain region and their corresponding behavioral dynamics.
@@ Appendix A Methodology details
Training Details of Neural LVM.  The neural signal dimensions for the brain regions are as follows: MOs_L: 14 dimensions, MOs_R: 14 dimensions, VIS_L: 24 dimensions, VIS_R: 21 dimensions, SSp_L: 23 dimensions, and SSp_R: 22 dimensions. Both the probabilistic encoder and decoder of the neural LVM are based on an RNN architecture Fabius and Van Amersfoort [2014]. Mean squared error (MSE) is used for both the neural reconstruction and behavior decoding loss. We use the Adam Optimizer Kingma and Ba [2014] for optimization and the learning rate is set as 0.001. The batch size is uniformly set to 32. The latent subspace factor number is fixed at 6, which is the same as the number of behaviors of interest. We employ the dropout technique Srivastava et al. [2014] and the ReLU activation function Rasamoelina et al. [2020] between layers in our probabilistic encoder and decoder neural networks.
Training Details of Video Diffusion Models.  We adopt the architecture of the VDM of 3D-UNet [Ho et al., 2022b] with the $\epsilon$-parameterization. We use both spatial attention and spatial convolutions. The temporal convolutions are used to maintain consistency between frames. The embedding input size to the UNet architecture is set as $32$ and the UNet has three downsampling and upsampling layers. The diffusion timestep is set as $200$. The training batch size is set as $64$, with a learning rate of 0.001. We use Group Normalization.
@@ Appendix B In-depth Investigation on the neural LVM module across brain regions
Metrics \ Method Session-1 Session-2
Standard VAE Ours Standard VAE Ours
@@ Appendix C Video Generation Results on Various Behaviors of Interests
Using Figure 9 as an example, for the naïve latent manipulation method, the generated frames are in a reasonable form. Nevertheless, the frame differences are still intertwined, and the latent factor of “Paw-(y)” heavily affects the “Spout” movement. Meanwhile, for classifier-free guidance, the trajectories focus on the mouse movements, but they are still entangled with the “Chest” movements. In contrast, the results of BeNeDiff show more specificity to the targeted behavior of interest. The inter-frame differences in BeNeDiff’s results are clearly specified to the “Paw-(y)” movements, and the temporal evolution of the neural dynamics is coherent with real-world mouse paw trajectories. The generated results in Figures 10 and 11 show a similar trend, demonstrating specificity to the Paw-(x)” and Spout” factors.
Figure 9: Generated Single-trial Behavioral Videos with Latent Factor Guidance from the bottom view. Compared to baseline methods, we observe that the neural dynamics of a latent factor in the results of BeNeDiff show specificity to the “Paw-(y)” movements.
Figure 10: Generated Single-trial Behavioral Videos with Latent Factor Guidance from the bottom view. Compared to baseline methods, we observe that the neural dynamics of a latent factor in the results of BeNeDiff show specificity to the “Paw-(x)” movements.
Figure 11: Generated Single-trial Behavioral Videos with Latent Factor Guidance from the bottom view. Compared to baseline methods, we observe that the neural dynamics of a latent factor in the results of BeNeDiff show specificity to the “Spout” movements.
Appendix D Learnt Neural Latent Trajectories of BeNeDiff across various brain regions
Figure 12: Learnt Neural Latent Trajectories of BeNeDiff across various brain regions. It is difficult to clearly visualize the specific motion encoded by each region and to distinguish how different the motions are encoded across brain regions.
Appendix E Video Generation Results on Various Brain Regions of the Left Hemisphere
Figure 13: Generated video frame differences across the left hemisphere regions. The red dots in the figure indicate paw appearances.
Appendix F Discussion and Limitation
Our study introduces BeNeDiff, a novel approach leveraging behavior-informed latent variable models and generative diffusion models to uncover and interpret neural dynamics. Through empirical validation, we demonstrate that BeNeDiff effectively identifies a disentangled neural subspace and synthesizes behavior videos that provide interpretable insights into neural activities associated with distinct behaviors of interest. However, for the neural latent variable model (LVM) module, there exists a balance between disentangling the neural subspace with behavior semantics and maintaining neural reconstruction performance. For each brain region and session, at this stage, a careful hyper-parameter search is necessary to balance the weight between these two components. For the generative video diffusion module, we implement the neural encoder (classifier for guidance) as a linear regressor for interpretability. This linear assumption can be relaxed later for improved guidance performance.
@@ 8. Answer: [Yes]
  9. Justification: In Section F Discussion.